If the rate of inflation is
3.4%
per year, the future price
pt
(in dollars) of a certain item can be modeled by the following exponential function, where t is the number of years from today.
=p(t)3000(1.034)
Find the current price of the item and the price 9 years from today.
Round your answers to the nearest dollar as necessary.

Answer by stanbon(75887)

f the rate of inflation is 3.4% per year, the future price
p(t) (in dollars) of a certain item can be modeled by the following exponential function, where t is the number of years from today.
p(t) = 3000(1.034)^t
Find the current price of the item and the price 9 years from today.
----
Current price = p(0) = $3000
-----
p(9) = 3000*1.034^9 = $4053.28
--------------
Cheers,
Stan H.
---------------